OpenStudy (anonymous):

please help. how can i determine the [OH−] of a 0.17M ammonia solution? that was all the information that was given. where do i begin?

OpenStudy (anonymous):

NH3 + H2O --> OH- + NH4+ 0.17 0 0 -x +x +x 0.17-x x x Kb of NH3 = [OH][NH4+] / [NH3] = x^2 / (0.17-x) We need to know the Kb of NH3

OpenStudy (anonymous):

So I looked up the Kb of NH3 and found it's about 1.82E-5. Since it's such a small number, we can assume "x" is negligible in size relative to the initial NH3 concentration of 0.17M thus... 1.82E-5 = x^2 / 0.17 --> x = [OH-] = 1.76E-3M

OpenStudy (anonymous):

So in questions like these, I would most likely have to look up kb/ka for the given base or acid? That's if it's not provided. And when in doubt, do that chart you did at the top?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem